A manufactor determines that the profit in dollars for manufactoring n units is P=2n^2-10n-175 (Assume that n is a positive integer). How many units are produced when the profit is $425?
Answer by stanbon(75887)

A manufactor determines that the profit in dollars for manufactoring n units is P=2n^2-10n-175 (Assume that n is a positive integer). How many units are produced when the profit is $425?
---------------------
2n^2-10n-175 = 425
----
2n^2-10n-600 = 0
---
n^2-5n-300 = 0
----
(n-20)(n+15) = 0
---
Positive solution:
n = 20 units
================
Cheers,
Stan H.